Three cards are drawn from a standard deck of cards, with replacement. True of false, and explain: (a) The chance of getting all hearts is 1.525% (b) The chance of getting at least one heart is 25% + 25% + 25% = 75%.
Answer #1

Saluboni Number of beasts = 13 Total number of cards=52 Tas PC of getting all hearts). . 13.) = 0.01562 à 11.562]. [False) (b
Answer #2

SOLUTION :


a.


P( 3 hearts from 3 draws with replacement)

= 13/52 * 13/52 * 13/52

= 1/64

= 0.015625

= 1.5625 %  . 

(It is same as given in the statement.)


So, the given statement is TRUE. (ANSWER) 


b.


P( 1 heart at least from 3 draws with replacement)

= 1 - P(no heart)

= 1 - (39/52 * 39/52 * 39/52)

= 1 - 27/64 

= 37/64

= 0.578125

= 57.8125%

 ( Statement tells that it is 75%)


So, the given statement is FALSE. (ANSWER)

answered by: Tulsiram Garg

> Part a : Please note that : the statement tells that it is 1.525 % but it should be 1.5625 %. So, the statement is FALSE.

Tulsiram Garg Fri, Dec 17, 2021 7:34 AM